A 505-turn solenoid has a radius of 7.20 mm and an overall length of 14.2 cm (a) What is its inductance? mH T I (b) If the solenoid is connected in series with a 2.50-92 resistor and a battery, what is the time constant of the circuit? ms
Answer #1

a) inductance = uN^2A/ l

= 4pi*10^-7*505^2 * pi*(0.0072)^2 / 0.142

= 3.6*10^-4 H = 0.36 mH

b) time constant = L/R

= 3.6*10^-4 / 2.5 = 1.47*10^-7 s